A spaceship's orbital maneuver requires a speed increase of 1.35  103 m/s. If its engine has an exhaust speed of 2.25  103 m/s, determine the required ratio Mi/Mf

 of its initial mass to its final mass. (The difference Mi − Mf

 equals the mass of the ejected fuel.)



____.



1
Expert's answer
2021-10-04T14:35:25-0400

According to Raw q conservation q momentum we have:

"(M+Δm)V = M(v+Δv)+Δm(v-v_e) \\

mΔv=v_eΔm \\

Δm=-ΔM \\

MΔv =-v_eΔM \\

v_f-v_i =v_eln(\frac{M_i}{M_f}) \\

\frac{v_f-v_i}{v_e} = ln(\frac{M_i}{M_f}) \\

\frac{1.35 \times 10^3}{2.25 \times 10^3} = ln(\frac{M_i}{M_f}) \\

\frac{M_i}{M_f} = e^{\frac{1.35}{2.25}} = e^{0.6}=1.822"